# MIntRec2.0: A Large-scale Benchmark Dataset for Multimodal Intent Recognition and Out-of-scope Detection in Conversations

### v2

@@ 2 Related Work
Benchmark Datasets. Intent recognition is a substantial task in natural language processing (NLP) and is supported by a numerous of benchmark datasets. These datasets can be broadly categorized into two branches. The first branch originates from task-oriented dialogues and includes datasets like ATIS (Tür et al., 2010), SNIPS (Coucke et al., 2018), CLINC150 (Larson et al., 2019), BANKING77 (Casanueva et al., 2020). Notably, CLINC150 incorporates out-of-scope data to test system robustness. SIMMC 2.0 (Kottur et al., 2021) is a multimodal dataset focusing on the shopping domain, but it lacks intent annotations. The second branch stems from open-ended dialogues, represented by multi-turn dialogue datasets such as DailyDialog (Li et al., 2017) and SWBD (Godfrey et al., 1992). However, these datasets primarily offer dialogue acts and may not be well-suited for applications requiring specific intent classes. In recent years, there has been a growing interest in multimodal language datasets for both single-turn (Zadeh et al., 2016; 2018b; Yu et al., 2020) and multi-turn dialogues (Busso et al., 2008; Poria et al., 2019). EMOTyDA (Saha et al., 2020) is another large-scale multimodal dataset for multi-turn dialogues, but it only includes coarse-grained dialogue acts. Some studies have also explored visual or multimodal intents using image modality (Jia et al., 2021; Kruk et al., 2019). MIntRec (Zhang et al., 2022a) stands as the first multimodal intent recognition dataset for open-ended dialogues. MIntRec2.0 significantly expands in scale from 2,224 to 15,040 utterances and is designed to handle both out-of-scope utterances and multi-turn dialogues. A comparison between MIntRec2.0 and other benchmark intent datasets is presented in Table 1.
@@
| ATIS (Tür et al., 2010) | 17 | 6,371 | ✓ | Single-turn | ✗ | ✗ | ✓ | ✗ | ✗ |
| Snips (Coucke et al., 2018) | 7 | 14,484 | ✓ | Single-turn | ✗ | ✗ | ✓ | ✗ | ✗ |
| CLINC150 (Larson et al., 2019) | 150 | 23,700 | ✓ | Single-turn | ✓ | ✗ | ✓ | ✗ | ✗ |
| MDID (Kruk et al., 2019) | 7 | 1,299 | ✗ | - | ✗ | ✗ | ✓ | ✓ | ✗ |
| Intentonomy (Jia et al., 2021) | 28 | 14,455 | ✗ | - | ✗ | ✗ | ✗ | ✓ | ✗ |
| MIntRec (Zhang et al., 2022a) | 20 | 2,224 | ✓ | Single-turn | ✗ | ✗ | ✓ | ✓ | ✓ |
| MIntRec2.0 | 30 | 15,040 | ✓ | Multi-turn | ✓ | ✓ | ✓ | ✓ | ✓ |
@@
| Achieve goals | Ask for opinions | Request others’ views or thoughts on a particular matter (e.g., asking for others’ perspectives). |
@@
Annotation Process. Six college students proficient in English are employed to perform multimodal label annotation. They are provided with a comprehensive guidebook detailing intent interpretations and application scenarios and are only permitted to begin after achieving high accuracy on seed examples. The annotators are evenly divided into two groups and assigned to annotate half of the data simultaneously. To facilitate their work, a user-friendly annotation platform with a unified database has been developed (Appendix I). Each worker is tasked with analyzing the speaker’s intention in a video segment by combining text, video, and audio information. They are then required to choose from a set of 30 known intent tags, as well as an OOS tag. The final label for each utterance is determined through majority voting, with at least two out of three votes required to reach a consensus. We operate under the assumption that each utterance has a single intent, and the rationale for not opting for multi-intent labeling is elaborated in Appendix J. To mitigate potential issues, utterances that receive three different votes are excluded from our dataset.
Annotation Results. We have successfully collected 1,245 high-quality dialogues to create the MIntRec2.0 dataset. This dataset consists of 9,304 in-scope and 5,736 out-of-scope utterances with multimodal labels. The statistics of the dataset are presented in Table 3. To assess annotation reliability, we calculate the Fleiss’s kappa statistics for each of our six annotators to measure interrater reliability. The Fleiss’s kappa scores range from 0.66 to 0.70, averaging 0.69. This indicates a level of substantial agreement, as defined in (McHugh, 2012). The distribution of the dataset across three different data sources is illustrated in Figure 3. Superstore, The Big Bang Theory, and Friends contribute 53%, 22%, and 25% of the dataset, respectively. Each data source contains between 54.5% and 67.9% of in-scope utterances. The intent distribution of in-scope utterances is depicted in Figure 3, demonstrating a common long-tailed distribution similar to real-world scenarios. As expected, some intents such as inform, explain, doubt, and complain are more prevalent in daily life, while others like warn, refuse, emphasize, and flaunt tend to occur less in specific occasions and scenes. To ensure adequate training, each intent class contains more than 90 samples.
@@ 4 Benchmark Framework
Training. Following multimodal fusion, we employ the multimodal representations for training. For in-scope samples , we perform classification on using the cross entropy loss , where is the number of training samples, and is the set of known intent labels. For out-of-scope samples , we apply the outlier exposure (OE) (Hendrycks et al., 2018) loss, denoted as , to distinguish them from the in-scope samples and enhance the model’s robustness and its generalization ability for out-of-scope samples. Specifically, we use a uniform distribution over the known classes as soft targets. The definitions for losses are as follows:
@@ 5 Experiments
Implementation Details. We partition our dataset into training, validation, and testing sets, maintaining an approximate ratio of 7:1:2 for both dialogues and utterances. (Further details are provided in Appendix L). For the text modality, we utilize BERT as a powerful backbone consisting of 24 transformer layers implemented in the Huggingface transformers library (Wolf et al., 2020), to extract features with the dimension of 1024. For the video modality, we employ well-trained checkpoints of Mask R-CNN from the MMDetection toolbox (Chen et al., 2019) to extract features with the dimension of 256. For the audio modality, we use the pre-trained model WavLM, implemented in (Wolf et al., 2020) to extract features with the dimension of 768. In single-turn dialogues, we apply zero-padding with a maximum sequence length of 50, 180, and 400 for text, video, and audio features, respectively. The number of training epochs is set to 40, and the training batch size is set to 16 for all baselines. We employ AdamW (Loshchilov & Hutter, 2019) for optimization, implement our approach using PyTorch 1.13.1, and conduct experiments on Tesla V100-SXM2-32GB GPUs. For all experiments, we report the results averaged over five runs, using random seeds ranging from 0 to 4 (Additional hyper-parameters details are available in Appendix M).
@@
Benchmark Baselines. As text is the predominant modality in conversational multimodal intent recognition (Zhang et al., 2022a), we establish a robust baseline by fine-tuning BERT, comparing its performance with two multimodal fusion methods: MAG-BERT and MulT. We evaluate these methods in both single-turn and multi-turn conversations, focusing on in-scope classification and out-of-scope detection. For single-turn conversations, we use only in-scope utterances for training. The out-of-scope utterances are included in the testing set and treated as a separate class, following (Lin & Xu, 2019; Zhang et al., 2023b). For multi-turn conversations, we consider both in-scope and out-of-scope samples at the dialogue-level during training, and all the baselines utilize the context information as described in section 4. We conduct additional baselines related to dialogue intent classification in NLP and out-of-distribution detection across different sources in Appendices N and O, respectively. Besides, we test the performance of ChatGPT on our dataset using both zero-shot and few-shot settings. In the zero-shot setting, ChatGPT is provided with the prompts of the label sets (e.g., 30 intent labels and one OOS) and an introduction to the task. In the few-shot setting, we use 10 dialogues with 227 utterances that cover all intent classes as the learning data (Details of the utilized prompts can be found in Appendix P). Finally, we invite ten evaluators to assess human performance. Each worker is assigned an equal portion of the testing set, ensuring they have not seen the data before. They receive the same background information as that provided to ChatGPT to ensure a fair comparison. Additionally, we provide them with more prior knowledge, consisting of 100 dialogues and 997 utterances, to explore human potential in addressing this complex problem. We average the predictions from all evaluators to obtain the final score.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62173195), the National Science and Technology Major Project towards the new generation of broadband wireless mobile communication networks of Jiangxi Province (03 and 5G Major Project of Jiangxi Province) (Grant No. 20232ABC03402), High-level Scientific and Technological Innovation Talents ”Double Hundred Plan” of Nanchang City in 2022 (Grant No. Hongke Zi (2022) 321-16), and supported by Hebei Natural Science Foundation (Grant No. F2022208006). We would like to thank Jiayan Teng, Zhaochen Yang, Shaojie Zhao, and Hao Li for their efforts during dataset construction.
@@ Appendix B Additional Related Work
Video Understanding. As a significant research field within computer vision, video understanding involves the extraction of valuable information from video content. Numerous methods have been developed to handle spatial and temporal data in videos, including the Two-Stream method, which comprises TDD (Wang et al., 2015), LRCN (Donahue et al., 2015), Fusion (Feichtenhofer et al., 2016), and TSN (Wang et al., 2016). This methodology integrates a secondary path to learn a video’s temporal information by training a convolutional neural network on the optical flow stream. However, these methods require extensive computation and storage capacity due to the pre-computation of optical flow.
@@
Intent Analysis. Intent analysis is an important research area in spoken language understanding (Qin et al., 2021). It plays a pivotal role in task-oriented dialogue systems, enabling the recognition of user queries’ intentions alongside the slot filling task (Wang et al., 2018b; Zhang et al., 2019). However, early research usually focus on the closed-world classification problem, lacking the capability to handle out-of-scope utterances encountered in real-world scenarios (Zhang et al., 2021a). To address this challenge, Lin & Xu (2019) first explore this task by employing margin loss to detect unknown intent. Zhang et al. (2021b) learn adaptive decision boundaries for each known class, thereby further reducing the open space risk. Yan et al. (2020) use Gaussian mixture models to tackle this problem and extends the task to zero-shot intent detection. Cheng et al. (2022) construct out-of-scope samples using manifold mixup technologies and employed soft labels for representation learning. Zhou et al. (2022) enhance intent representations to balance both empirical and open space risks with the aid of contrastive learning in the K-nearest neighbors space.
@@ Appendix C Performance of DialogueRNN
To leverage context information, existing methods typically use multimodal fusion representations to directly model the temporal information of contexts. However, we find this approach to be ineffective for our task. Specifically, we select DialogueRNN (Majumder et al., 2019), a method specifically designed for multimodal emotion detection in conversations, for evaluation. We conduct experiments under two settings: +1 and Outlier Exposure. The former treats the out-of-scope class as the (+1) class and trains using both intent classes and one out-of-scope class, while the latter employs the outlier exposure loss on out-of-scope data during training.
@@ Appendix H Application of Intent Labels
| Criticize | Point out and emphasize someone’s mistakes (e.g., yelling out someone’s problems). | ||
@@ Appendix I Multimodal Intent Annotation Platform
We partition our dataset into training, validation, and testing sets at an approximate ratio of 7:1:1 for both utterances and dialogues. Detailed statistics for each set, encompassing both in-scope and out-of-scope data, are presented in Table 10.
| Item | # Dialogues | # Utterances | # In-scope Utterances | # Out-of-scope Utterances |
@@
The results show that even state-of-the-art methods for open intent detection generally underperform compared to the BERT text classifier across most metrics. However, they do excel in identifying out-of-scope utterances, typically achieving higher F1-OOS scores. Notably, KNNCL also scores higher in accuracy.
@@
ChatGPT-10 Prompts: Here is a list of multiple multi-turn conversations. Each dictionary in the list represents a conversation paragraph, where each key-value pair represents an intent example as a key and its corresponding label as a value. Next time I will enter my request, please only reply ”received”. This is a list of given intent labels: [ Acknowledge, Advise, Agree, Apologise, Arrange, Ask for help, Asking for opinions, Care, Comfort, Complain, Confirm, Criticize, Doubt, Emphasize, Explain, Flaunt, Greet, Inform, Introduce, Invite, Joke, Leave, Oppose, Plan, Praise, Prevent, Refuse, Taunt, Thank, Warn, OOS], where OOS represents an unknown intent that is not intended otherwise. Now, you need to learn from the conversations that you were given in the last Q&A, and then I’ll provide you with a dialog that contains utterances in it, and these utterances are given in order and can be considered as contextual. Now, for each utterance that requires you to use the knowledge you gained from the given conversations, select a label as output from the given list of labels: for the following given dialog, in this format: Original sample: Intent labels output.
@@
Limitations: This study presents several limitations that warrant acknowledgment. First, deploying this system in real-world settings necessitates collecting personal data, including facial expressions, voice, and text, thereby raising critical privacy concerns requiring meticulous attention. Second, the issue of liability remains ambiguous, especially in sensitive applications such as medical diagnosis, should the technology produce erroneous results. Third, our training dataset may lack comprehensive representation across diverse cultural backgrounds, potentially resulting in misunderstandings or the perpetuation of stereotypes. Lastly, substantial opportunities exist for enhancing the system’s performance, particularly in effectively utilizing context information and out-of-scope sample data and incorporating non-verbal modalities.

### v3

@@ 2 Related Work
Benchmark Datasets. Intent recognition is a substantial task in natural language processing (NLP) and is supported by a numerous of benchmark datasets. These datasets can be broadly categorized into two branches. The first branch originates from task-oriented dialogues and includes datasets like ATIS (TĂŒr et al., 2010), SNIPS (Coucke et al., 2018), CLINC150 (Larson et al., 2019), BANKING77 (Casanueva et al., 2020). Notably, CLINC150 incorporates out-of-scope data to test system robustness. SIMMC 2.0 (Kottur et al., 2021) is a multimodal dataset focusing on the shopping domain, but it lacks intent annotations. The second branch stems from open-ended dialogues, represented by multi-turn dialogue datasets such as DailyDialog (Li et al., 2017) and SWBD (Godfrey et al., 1992). However, these datasets primarily offer dialogue acts and may not be well-suited for applications requiring specific intent classes. In recent years, there has been a growing interest in multimodal language datasets for both single-turn (Zadeh et al., 2016; 2018b; Yu et al., 2020) and multi-turn dialogues (Busso et al., 2008; Poria et al., 2019). EMOTyDA (Saha et al., 2020) is another large-scale multimodal dataset for multi-turn dialogues, but it only includes coarse-grained dialogue acts. Some studies have also explored visual or multimodal intents using image modality (Jia et al., 2021; Kruk et al., 2019). MIntRec (Zhang et al., 2022a) stands as the first multimodal intent recognition dataset for open-ended dialogues. MIntRec 2.0 significantly expands in scale from 2,224 to 15,040 utterances and is designed to handle both out-of-scope utterances and multi-turn dialogues. A comparison between MIntRec 2.0 and other benchmark intent datasets is presented in Table 1.
@@
| ATIS (TĂŒr et al., 2010) | 17 | 6,371 | â | Single-turn | â | â | â | â | â |
| Snips (Coucke et al., 2018) | 7 | 14,484 | â | Single-turn | â | â | â | â | â |
| CLINC150 (Larson et al., 2019) | 150 | 23,700 | â | Single-turn | â | â | â | â | â |
| MDID (Kruk et al., 2019) | 7 | 1,299 | â | - | â | â | â | â | â |
| Intentonomy (Jia et al., 2021) | 28 | 14,455 | â | - | â | â | â | â | â |
| MIntRec (Zhang et al., 2022a) | 20 | 2,224 | â | Single-turn | â | â | â | â | â |
| MIntRec 2.0 | 30 | 15,040 | â | Multi-turn | â | â | â | â | â |
@@
| Achieve goals | Ask for opinions | Request othersâ views or thoughts on a particular matter (e.g., asking for othersâ perspectives). |
@@
| # data sources | 3 |
| # intents classes | 30 |
| # dialogues | 1,245 |
| # utterances | 15,040 |
| # in-scope utterances | 9,304 |
| # out-of-scope utterances | 5,736 |
| # words in utterances | 118,477 |
| # unique words in utterances | 9,524 |
| Average length of utterances | 7.9 |
| Maximum length of utterances | 46 |
| Average video clip duration | 3.0 (s) |
| Maximum video clip duration | 19.9 (s) |
| Video hours | 12.3 (h) |
@@
Annotation Process. Six college students proficient in English are employed to perform multimodal label annotation. They are provided with a comprehensive guidebook detailing intent interpretations and application scenarios and are only permitted to begin after achieving high accuracy on seed examples. The annotators are evenly divided into two groups and assigned to annotate half of the data simultaneously. To facilitate their work, a user-friendly annotation platform with a unified database has been developed (Appendix I). Each worker is tasked with analyzing the speakerâs intention in a video segment by combining text, video, and audio information. They are then required to choose from a set of 30 known intent tags, as well as an OOS tag. The final label for each utterance is determined through majority voting, with at least two out of three votes required to reach a consensus. We operate under the assumption that each utterance has a single intent, and the rationale for not opting for multi-intent labeling is elaborated in Appendix J. To mitigate potential issues, utterances that receive three different votes are excluded from our dataset.
Annotation Results. We have successfully collected 1,245 high-quality dialogues to create the MIntRec 2.0 dataset. This dataset consists of 9,304 in-scope and 5,736 out-of-scope utterances with multimodal labels. The statistics of the dataset are presented in Table 3. To assess annotation reliability, we calculate the Fleissâs kappa statistics for each of our six annotators to measure interrater reliability. The Fleissâs kappa scores range from 0.66 to 0.70, averaging 0.69. This indicates a level of substantial agreement, as defined in (McHugh, 2012). The distribution of the dataset across three different data sources is illustrated in Figure 3. Superstore, The Big Bang Theory, and Friends contribute 53%, 22%, and 25% of the dataset, respectively. Each data source contains between 54.5% and 67.9% of in-scope utterances. The intent distribution of in-scope utterances is depicted in Figure 3, demonstrating a common long-tailed distribution similar to real-world scenarios. As expected, some intents such as inform, explain, doubt, and complain are more prevalent in daily life, while others like warn, refuse, emphasize, and flaunt tend to occur less in specific occasions and scenes. To ensure adequate training, each intent class contains more than 90 samples.
@@ 4 Benchmark Framework
Training. Following multimodal fusion, we employ the multimodal representations for training. For in-scope samples , we perform classification on using the cross entropy loss , where is the number of training samples, and is the set of known intent labels. For out-of-scope samples , we apply the outlier exposure (OE) (Hendrycks et al., 2018) loss, denoted as , to distinguish them from the in-scope samples and enhance the modelâs robustness and its generalization ability for out-of-scope samples. Specifically, we use a uniform distribution over the known classes as soft targets. The definitions for losses are as follows:
@@ 5 Experiments
Implementation Details. We partition our dataset into training, validation, and testing sets, maintaining an approximate ratio of 7:1:2 for both dialogues and utterances. (Further details are provided in Appendix L). For the text modality, we utilize BERTLARGE as a powerful backbone consisting of 24 transformer layers implemented in the Huggingface transformers library (Wolf et al., 2020), to extract features with the dimension of 1024. For the video modality, we employ well-trained checkpoints of Mask R-CNN from the MMDetection toolbox (Chen et al., 2019) to extract features with the dimension of 256. For the audio modality, we use the pre-trained model WavLM, implemented in (Wolf et al., 2020) to extract features with the dimension of 768. In single-turn dialogues, we apply zero-padding with a maximum sequence length of 50, 180, and 400 for text, video, and audio features, respectively. The number of training epochs is set to 40, and the training batch size is set to 16 for all baselines. We employ AdamW (Loshchilov & Hutter, 2019) for optimization, implement our approach using PyTorch 1.13.1, and conduct experiments on Tesla V100-SXM2-32GB GPUs. For all experiments, we report the results averaged over five runs, using random seeds ranging from 0 to 4 (Additional hyper-parameters details are available in Appendix M).
@@
Benchmark Baselines. As text is the predominant modality in conversational multimodal intent recognition (Zhang et al., 2022a), we establish a robust baseline by fine-tuning BERTLARGE, comparing its performance with two multimodal fusion methods: MAG-BERT and MulT. We evaluate these methods in both single-turn and multi-turn conversations, focusing on in-scope classification and out-of-scope detection. For single-turn conversations, we use only in-scope utterances for training. The out-of-scope utterances are included in the testing set and treated as a separate class, following (Lin & Xu, 2019; Zhang et al., 2023b). For multi-turn conversations, we consider both in-scope and out-of-scope samples at the dialogue-level during training, and all the baselines utilize the context information as described in section 4. We conduct additional baselines related to dialogue intent classification in NLP and out-of-distribution detection across different sources in Appendices N and O, respectively. Besides, we test the performance of ChatGPT on our dataset using both zero-shot and few-shot settings. In the zero-shot setting, ChatGPT is provided with the prompts of the label sets (e.g., 30 intent labels and one OOS) and an introduction to the task. In the few-shot setting, we use 10 dialogues with 227 utterances that cover all intent classes as the learning data (Details of the utilized prompts can be found in Appendix P). Finally, we invite ten evaluators to assess human performance. Each worker is assigned an equal portion of the testing set, ensuring they have not seen the data before. They receive the same background information as that provided to ChatGPT to ensure a fair comparison. Additionally, we provide them with more prior knowledge, consisting of 100 dialogues and 997 utterances, to explore human potential in addressing this complex problem. We average the predictions from all evaluators to obtain the final score.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62173195), the National Science and Technology Major Project towards the new generation of broadband wireless mobile communication networks of Jiangxi Province (03 and 5G Major Project of Jiangxi Province) (Grant No. 20232ABC03402), High-level Scientific and Technological Innovation Talents âDouble Hundred Planâ of Nanchang City in 2022 (Grant No. Hongke Zi (2022) 321-16), and supported by Hebei Natural Science Foundation (Grant No. F2022208006). We would like to thank Jiayan Teng, Zhaochen Yang, Shaojie Zhao, and Hao Li for their efforts during dataset construction.
@@ Appendix B Additional Related Work
Video Understanding. As a significant research field within computer vision, video understanding involves the extraction of valuable information from video content. Numerous methods have been developed to handle spatial and temporal data in videos, including the Two-Stream method, which comprises TDD (Wang et al., 2015), LRCN (Donahue et al., 2015), Fusion (Feichtenhofer et al., 2016), and TSN (Wang et al., 2016). This methodology integrates a secondary path to learn a videoâs temporal information by training a convolutional neural network on the optical flow stream. However, these methods require extensive computation and storage capacity due to the pre-computation of optical flow.
@@
Intent Analysis. Intent analysis is an important research area in spoken language understanding (Qin et al., 2021). It plays a pivotal role in task-oriented dialogue systems, enabling the recognition of user queriesâ intentions alongside the slot filling task (Wang et al., 2018b; Zhang et al., 2019). However, early research usually focus on the closed-world classification problem, lacking the capability to handle out-of-scope utterances encountered in real-world scenarios (Zhang et al., 2021a). To address this challenge, Lin & Xu (2019) first explore this task by employing margin loss to detect unknown intent. Zhang et al. (2021b) learn adaptive decision boundaries for each known class, thereby further reducing the open space risk. Yan et al. (2020) use Gaussian mixture models to tackle this problem and extends the task to zero-shot intent detection. Cheng et al. (2022) construct out-of-scope samples using manifold mixup technologies and employed soft labels for representation learning. Zhou et al. (2022) enhance intent representations to balance both empirical and open space risks with the aid of contrastive learning in the K-nearest neighbors space.
@@ Appendix C Performance of DialogueRNN
To leverage context information, existing methods typically use multimodal fusion representations to directly model the temporal information of contexts. However, we find this approach to be ineffective for our task. Specifically, we select DialogueRNN (Majumder et al., 2019), a method specifically designed for multimodal emotion detection in conversations, for evaluation. We conduct experiments under two settings: +1 and Outlier Exposure. The former treats the out-of-scope class as the (+1)th class and trains using both intent classes and one out-of-scope class, while the latter employs the outlier exposure loss on out-of-scope data during training.
@@ Appendix H Application of Intent Labels
| Criticize | Point out and emphasize someoneâs mistakes (e.g., yelling out someoneâs problems). | ||
@@ Appendix I Multimodal Intent Annotation Platform
We partition our dataset into training, validation, and testing sets at an approximate ratio of 7:1:1 for both utterances and dialogues. Detailed statistics for each set, encompassing both in-scope and out-of-scope data, are presented in Table  10.
| Item | #D | #U | # U (In-scope) | #U (Out-of-scope) |
@@
The results show that even state-of-the-art methods for open intent detection generally underperform compared to the BERTLARGE text classifier across most metrics. However, they do excel in identifying out-of-scope utterances, typically achieving higher F1-OOS scores. Notably, KNNCL also scores higher in accuracy.
@@
ChatGPT-10 Prompts: Here is a list of multiple multi-turn conversations. Each dictionary in the list represents a conversation paragraph, where each key-value pair represents an intent example as a key and its corresponding label as a value. Next time I will enter my request, please only reply âreceivedâ. This is a list of given intent labels: [ Acknowledge, Advise, Agree, Apologise, Arrange, Ask for help, Asking for opinions, Care, Comfort, Complain, Confirm, Criticize, Doubt, Emphasize, Explain, Flaunt, Greet, Inform, Introduce, Invite, Joke, Leave, Oppose, Plan, Praise, Prevent, Refuse, Taunt, Thank, Warn, OOS], where OOS represents an unknown intent that is not intended otherwise. Now, you need to learn from the conversations that you were given in the last Q&A, and then Iâll provide you with a dialog that contains utterances in it, and these utterances are given in order and can be considered as contextual. Now, for each utterance that requires you to use the knowledge you gained from the given conversations, select a label as output from the given list of labels: for the following given dialog, in this format: Original sample: Intent labels output.
@@
Limitations: This study presents several limitations that warrant acknowledgment. First, deploying this system in real-world settings necessitates collecting personal data, including facial expressions, voice, and text, thereby raising critical privacy concerns requiring meticulous attention. Second, the issue of liability remains ambiguous, especially in sensitive applications such as medical diagnosis, should the technology produce erroneous results. Third, our training dataset may lack comprehensive representation across diverse cultural backgrounds, potentially resulting in misunderstandings or the perpetuation of stereotypes. Lastly, substantial opportunities exist for enhancing the systemâs performance, particularly in effectively utilizing context information and out-of-scope sample data and incorporating non-verbal modalities.